OpenStudy (anonymous):

determine the equation of the tangent to the curve defined by y=xe^-x at the point A(1,e^-1)

OpenStudy (anonymous):

yea, thats what i got, but the answer is y=1/e

OpenStudy (anonymous):

now i believe i gotta substitute the coordinates to calculate b value

OpenStudy (anonymous):

with the given points

OpenStudy (anonymous):

slope at point A is 0. so your equation is: y-1/e =0

OpenStudy (anonymous):

same thing...

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

, prove that y"=y/25